OpenStudy (anonymous):

solve the equation by factoring 3x(x-4)=7x^2-13x

OpenStudy (anonymous):

start with \[3 x^2-12 x = 7 x^2-13 x\] then \[x-4 x^2 = 0\]

OpenStudy (anonymous):

or if you prefer \[4x^2-x=0\] \[x(4x-1)=0\] \[x=0, x=\frac{1}{4}\]

OpenStudy (anonymous):

How did you get that answer?